OpenStudy (anonymous):

Gold has a density of 19.3 g/cm^3.The gold in a ring has a mass of 15 g. Work out the volume of the gold in the ring.

OpenStudy (anonymous):

Density = mass/volume... Just plug in the values.

OpenStudy (anonymous):

d=m/v v=m/d 15/19.3

OpenStudy (anonymous):

=0.778 cm^3